what is the slope intercept form?
3x+7=0

Answer by stanbon(75887)

what is the slope intercept form?
3x+7=0
--------------
The slope-intercept form is y = mx+b where m is the slope and b is the intercept.
-----------------
Your problem has no y term.
Its equation is 3x = -7
x = -7/3 which is a vertical line with undefined slope
and no y-intercept
==========================
Cheers,
Stan H.
PS: Send reply to stanbon@comcast.net as the algebra.com
feedback program is not working.